# Classification of Histopathology Slides with Persistence Homology Convolutions

### v1

@@ 2.1 Persistent Homology
    These intervals correspond to holes (voids) in $X.$ $b$ and $d$ are the maximum and minimum $y$-coordinates of the hole, respectively.
On the right of Figure 3 is a visualization of the diagram generated from the example to its left. The set $X$ has one component and two holes, so it will have one intervals of Type 2 and two intervals of Type 3. We describe these and the other intervals in the diagram in order, scanning from below.
Scanning upwards, observe that a Type 2 feature appear in the structure at $t=0$ continuing till $t=5$. This feature is recorded in the diagram by the point $a_{2}$. Additionally, in between this sweep a Type 1 feature appears. At $t=0.5,$ the bottom left leg appears in the frame as a separate component. This leg eventually merges with the larger structure at $t=2$ and is recorded in the diagram as the point $a_{1}$. These measurements describe 00-dimensional features ($\mathrm{Ext}_{0}(f)$) and capture information regarding the connectedness of the structure and the length of each component.
Upon descending the structure Type 3 features begin to emerge. At $t=4.5$ the top right leg bounds the region in red below the horizontal line which persists until $t=3$ and is recorded in the diagram by the point $a_{5}$. This is similarly repeated for features $a_{3}$ and $a_{4}$. These are the $1$-dimensional features of the data ($\mathrm{Ext}_{1}(f)$) and are used to characterize the length of the bounded regions (which we associate with a cell). Larger values of $p$ are used to determine higher-dimensional topological features but are not present for two dimensional image data. In our study we will restrict our focus to the case when $p=1$ and exclude summaries generated when $p=0$ and intend to reincorporate them in future work. We hypothesize that the summary $\mathrm{Ext}_{1}(f)$ is most important in distinguishing between classes of tumor growth—we expand on how this is computationally implemented in Section 4.1.2.
@@ 2.2 Vectorization of Persistence Data
As described above, the PHC data consists of collections of sets of intervals, one for each window. This incompatible with the machine learning algorithms we intend to use, which take vectors as input. In particular, note that the number of intervals may vary between windows. As such, we convert the PDs into vectors in a fixed-dimensional Euclidean space. There are multiple vectorization methods for persistence diagrams [35]. A precise mathematical definition of the vectorization operation is found in [41].
There are multiple vectorization methods for persistence diagrams. In our study, we utilize the notion of a persistence image developed by Adams et al [7]. This representation converts a diagram into a finite-dimensional vector in $\mathbb{R}^{n\times n}$ — denoted $\mathbf{I}_{n\times n}(z)$ — that may be interpreted as a discretized heat map of the scatter plot of interval endpoints of $\text{Ext}_{1}(f)$. See [7] for a complete definition. The persistence image is stable with respect to input noise. In one study, it was shown to perform better than alternative vectorizations in machine learning applications [7].
@@ 3 Persistent Homology Convolutions
$[X\star k](x,y)=\sum_{\tau_{x}\in\mathbb{Z}}{\sum_{\tau_{y}\in\mathbb{Z}}{k(x,y)X(x-\tau_{x},y-\tau_{y})}}.$
Intuitively, the kernel $k(x,y)$ is used to extract relevant features from $f(x,y)$ by performing computations with the kernel as it translates over the entire image. The output of this operation is a feature map that contains synthesized information regarding the original image and is then subsampled (pooled) and is then (generally) repeated multiple times before being feed to the multilayer perception. The convolution operator is translation invariant but the process of convolving and pooling does not necessarily preserve the topology in an image. To address these shortcomings, we present Persistent Homology Convolutions. This operation shares similar attributes to the classic convolution operator, but instead measures local topological features while maintaining that the operator is translation invariant to such measurements. We present this idea on an $N\times N$ grayscale image $X\in\mathcal{G}_{N}$ where $\mathcal{G}_{N}$ denotes the set $[0,1]^{[[N]]\times[[N]]}$ with $[[N]]=\{0,1,\dots,N\}$.
@@
Let $X\in\mathcal{G}_{N}$ be an $N\times N$ grayscale image. For $M<N$ consider the function $T_{(x,y)}\colon\mathcal{G}_{N}\to\mathcal{G}_{M}$ which maps $X$ to an $M\times M$ grayscaled image by translating the $(x,y)$ coordinate of $X$ to the origin and restricting computations to $[[M]]^{2}$. Given a kernel $k(x,y)$ and a function $\mathcal{F}$ assigning a filtration to an $M\times M$ subimage, we define the Persistent Homology Convolution operator as
$[X\star_{PH}k]=\sum_{i=1}^{K}{\sum_{j=1}^{K}{k(c\cdot i,c\cdot j)}\mathbf{vec}(\mathrm{PH}_{p}(\mathcal{F}(T_{(c\cdot i,c\cdot j)}(X)))}$
@@
Here, $c$ acts as the stride length used to translate the window. To prevent the window from being translated outside of the boundary of the image, the $x$-coordinate and $y$-coordinate translations are bounded between $1\leq i,j\leq K$. In our study, the filtration $\mathcal{F}(T_{(c\cdot i,c\cdot j)}(X))$ corresponds to the extended sublevel-set filtration of the height function on the $M\times M$ subimage of $X$ obtained by thresholding (as described in the next section). Since persistence is computed with respect to dimension one height PH, the final term is expressed as $\mathbf{I}_{n\times n}(\mathrm{Ext}_{1}(f(T_{(32\cdot i,32\cdot j)}(X))))$ where the stride was set to 32 (see Section 4.2). Similarly, the methods used by both [3] and [30] to compute persistence on histology slides can be presented as $\mathrm{PH}_{p}(\mathrm{St}_{\underline{\hskip 5.69054pt}}(X))$ where $\mathrm{St}_{\underline{\hskip 5.69054pt}}(X)$ represents the lower-star filtration on the magnitude of pixels in $X$ and $p=0,1$. The map $T_{(i,j)}$ is removed from the expression since the computations are performed globally on $X$ rather than on windows.
For the purpose of this study, the kernel is a linear operator (an arbitrary $q\times q$ matrix) whose entries are optimized via backpropagation during model training. The kernel performs feature extraction on the vectorized data to optimally search for the relevant topological signatures which it uses to distinguish between the tumor classes. We hope to use a similar approach in the future to adaptively optimize a weighting function in the vectorization operation.
@@ 4.1 Experimental Setup
We study three distinct representations of a histopathology slide: a grayscaled, thresholded image, global persistent homology computed for the entire image, and local persistent homology in the form of PHCs. Our objective is to compare the performance of models trained with each of these three representations with respect to multiple performance metrics. We also compare model performance of PHCs based on three different filtrations.
@@ 4.1.1 Histopathology Dataset
We apply CNNs to a dataset of histopathology slides for the diagnosis of Osteosarcoma (Ost.), a rare form of bone cancer. It is available at the Cancer Imaging Archive [46, 4]. The dataset consists of 1144 RGB images with resolution (1024, 1024, 3). The images are separated into three classes: non-tumorous (47%), necrotic tumo (23%), and viable tumor (30%). It is desirable that the dataset is balanced in each class with roughly 381 images (roughly a third of the total number of images). To achieve this, we resample 381 non-tumorous images and perform image augmentation (rotations and reflections) on the remaining classes to balance them. Since homology is invariant to rotations and reflections, the global PH data will contain multiple identical summaries. This illustrates that certain forms of data augmentation cannot be used when persistence is computed globally without introducing redundancies. The resulting experimental dataset contains 1143 RGB images and serves as our base-line comparison against the PH-based data.
@@ 4.1.2 PHC Data Generation
There are several preprocessing steps that are taken to condition the image data before computing persistent homology. These standard methods are applied to enhance the geometric features in the tissue structure, enabling more accurate and robust summaries. Note that we apply the same preprocessing to the data used for each of the three methods (that is: global persistence, local persistence, and image data) to ensure a fair comparison. The basic procedure is as follows:
@@
    Thresholding and image erosion are utilized to emphasize the tissue architecture.
@@
    All pixels remaining after the thresholding (the boundary of the cells) are appended to a data structure known as a simplex tree along with edges from neighboring pixels whose values are nonzero.
@@
Figure 4: Our preprocessing pipeline for image classification using PHC. As the convolution operator slides across the image several sub-tasks are performed. The sub-image is first thresholded before being used for complex creation. Next, the persistence of the complex is computed and vectorized. The data is collated and used for model training.
@@
Thresholding is used to emphasize the boundary of the cells and remove noise. We experimented with multiple threshold values ranging from 160-200 and found $k=200$ was the most optimal. Dilation is applied once to “thicken” the boundary of the cells and fill in any small, “noisy” holes using a $2\times 2$ kernel.
To compute persistence, the conditioned image data from Step 2 must be converted into a data structure known as simplex tree, a trie data structure used to efficiently represent any general simplicial complex [33]. We create a simplex tree by adding edges between adjacent pixels in the thresholded set; this is called the adjacency complex of the image. There are other methods of creating a simplicial complex from the image data such as creating the alpha complex on the pixels of a thresholded image or computing lower-star filtration on the magnitude of the pixels.
@@
so $n=20$ was fixed. The output PHC data of a single image is a 3-dimensional tensor of the form $(256,20,20)$ which corresponds to $256$ persistence images (one at each locality) of resolution $20\times 20$.
All persistence-based computations are performed using the computational topology library GUDHI [32]. Our pipeline is visually summarized in Figure 4. In its current form, most of the PHC data generation is a separate preprocessing step applied on the image data outside of the training cycle except for the optimizing the weights of $k(x,y)$. Further optimization can be performed on the parameter space which we plan to expand upon in future work, but is outside the scope of this study.
@@ 4.2 Model, Training, and Implementation Details
We describe the CNN architecture used for our computational study. Further, our study indicates that a small CNN is sufficient to achieve high performance, consisting only of two convolution and pooling layers (which update the weights of $k(x,y)$) followed by three dense layers for the multilayer perceptron. These models are small enough that most modern laptops have sufficient hardware to train them.
All models are created and trained using TensorFlow [36]. The exact specifications of the architecture is as follows (see [42] for an explanation of the terms used below). The kernel size is fixed to $(3,3)$ for both convolutions layers with a stride of two. ReLu is the default activation function used for all layers of the CNN except the last which uses the SoftMax function. The weights of the model are initialized using He Normal Initialization [16]. We employ Adam[43] for optimization with a fixed learning rate of $\alpha=0.001$. The loss functions used is categorical cross entropy given by the formula
@@
Similarly the convolution filter sizes vary from $\{8,16,32\}$ and layer size from $\{64,128,256,512,1028\}$. All of these hyperparameters were chosen using a Bayesian hyperparameter sweep implemented using Weights and Biases [18] to maximize for accuracy.
In addition to comparing model performance between the three distinct representations, we also test the combination of image data with persistence (images + global PH or images + local PH). To train a model with the combined representations of image and PH data the CNN model architecture is slightly modified. Instead, it includes two separate feature extraction blocks with the same number of convolution and pooling layers as before using the same sweep configuration listed above. The image and PH data are fed through each separate block and concatenated together before being fed through the dense layers. We hypothesize that the inductive biases (translational invariance, locality, and hierarchical learning [17]) that are built into CNNs help distinguish between topological signatures in the persistence image. All five distinct training sets were trained over a 1000 model initializations.
@@ 4.3 Evaluation Metrics
Accuracy=Correct ClassificationsAll $\textbf{Accuracy}=\frac{\text{Correct Classifications}}{\text{All Classifications}}$

### v3

@@ 2.1.3 Ordinary Persistent Homology
[…]
Larger values of $p$ correspond to higher-dimensional topological features which do not occur in this setting. Each type of feature captured Proposition 2 has the following biological interpretations.
Type 2 features correspond to various cellular and subcellular structures that are darker than their neighbors; the birth and death values respectively encode information about the tissue type of the structure and its neighbors. Type 2 features for the osteosarcoma dataset indicate the presence of a cell, nuclei, nuclear fragmentation (karyorrhexis), or other distinct tissue matter. The darkest structures in this histopathology data set are nuclei, and multinucleation events are recorded by Type 2 features with low death times (as shown in the second row of Figure 4). Type 3 features correspond analogously to structures that are lighter than their neighbors, such as empty voids between cells as illustrated in the third row of Figure 4). Intervals of both types of features whose corresponding critical points are in the interior of $D$ represent structures contained in the interior of $D,$ whereas the others encode information about cellular and subcellular structures adjacent to the boundary.
2.1.4 Extended Persistent Homology
[…]
Proposition 3.
[…]
  • Type 4.
[…]
  • Type 5.
[…]
@@ 2.2 Vectorization of Persistence Data
As described above, the PHC data consists of collections of sets of intervals, one for each window. This incompatible with the machine learning algorithms we intend to use, which take vectors as input. In particular, note that the number of intervals may vary between windows. As such, we convert the PDs into vectors in a fixed-dimensional Euclidean space. There are multiple vectorization methods for persistence diagrams (Ali et al., 2023). A precise mathematical definition of the vectorization operation is found in (Carrière et al., 2020).
There are multiple vectorization methods for persistence diagrams. In our study, we utilize the notion of a persistence image developed by (Adams et al., 2017). This representation converts a diagram into a finite-dimensional vector in $\mathbb{R}^{n\times n}$ — denoted $\mathbf{I}_{n\times n}(z)$ — that may be interpreted as a discretized heat map of the scatter plot of interval endpoints of $\text{XPH}_{1}(f)$. See (Adams et al., 2017) for a complete definition. The persistence image is stable with respect to input noise. In one study, it was shown to perform better than alternative vectorizations in machine learning applications (Adams et al., 2017).
@@ 3 Persistent Homology Convolutions
$k\star X](x,y)=\sum_{\tau_{x}\in\mathbb{Z}}{\sum_{\tau_{y}\in\mathbb{Z}}{k(x,y)X(x-\tau_{x},y-\tau_{y})}}.$ (1)
Intuitively, the kernel $k(x,y)$ is used to extract relevant features from $f(x,y)$ by performing computations with the kernel as it translates over the entire image. The output of this operation is a feature map that contains synthesized information regarding the original image and is then subsampled (pooled) and is then (generally) repeated multiple times before being feed to the multilayer perception. The convolution operator is translation invariant but the process of convolving and pooling does not necessarily preserve the topology in an image. To address these shortcomings, we present Persistent Homology Convolutions. This operation shares similar attributes to the classic convolution operator, but instead measures local topological features while maintaining that the operator is translation equivariant to such measurements. We present this idea on an $N\times N$ grayscale image $X\in\mathcal{G}_{N}$ where $\mathcal{G}_{N}$ denotes the set $0,1]^{[[N]]\times[[N]]}$ with $N]]=\{0,1,\dots,N\}$.
@@
Let $X\in\mathcal{G}_{N}$ be an $N\times N$ grayscale image. For $M<N$ consider the function $T_{(x,y)}\colon\mathcal{G}_{N}\to\mathcal{G}_{M}$ which maps $X$ to an $M\times M$ grayscaled image by translating the $(x,y)$ coordinate of $X$ to the origin and restricting computations to $M]]^{2}$. Given a kernel matrix $\mathbf{K}$ and a function $\mathcal{F}$ assigning a filtration to an $M\times M$ subimage, we define the Persistent Homology Convolution operator as
$\mathbf{K}\star_{PH}X]=\sum_{i=1}^{K}{\sum_{j=1}^{K}{\mathbf{K}\star}\mathbf{vec}(\mathrm{PH}_{p}(\mathcal{F}(T_{(c\cdot i,c\cdot j)}(X)))}$ (2)
@@
Here, $c$ acts as the stride length used to translate the window. To prevent the window from being translated outside of the boundary of the image, the $x$-coordinate and $y$-coordinate translations are bounded between $1\leq i,j\leq K$. In our study, the filtration $\mathcal{F}(T_{(c\cdot i,c\cdot j)}(X))$ corresponds to either the alpha complex, or a modified version of the extended lower star function on the $M\times M$ subimage of $X$ (as described in the next section). In our study, we focus on persistence computed in dimension one as it yielded the best results. The localized PH computation is expressed as $Y=\mathbf{I}_{n\times n}(\mathrm{XPH}_{1}(f(T_{(32\cdot i,32\cdot j)}(X))))$ where the stride was set to 32 (see Section 4.3). Similarly, the methods used by both (Qaiser et al., 2019) and (Lawson et al., 2019) to compute persistence on histology slides can be presented as $\mathrm{PH}_{p}(\mathrm{St}_{\underline{\hskip 5.69054pt}}(X))$ where $\mathrm{St}_{\underline{\hskip 5.69054pt}}(X)$ represents the lower star filtration on the magnitude of pixels in $X$ and $p=0,1$. The map $T_{(i,j)}$ is removed from the expression since the computations are performed globally on $X$ rather than on windows.
Using persistence images, the output local PH computation is a topological feature vector in $Y\in\mathbb{R}^{n\times n}$ (as stated above). The $n\times n$ array is then convolved with the kernel matrix $\mathbf{K}$ whose entries are $\mathbf{K_{ij}}=k(w,T_{\left(k,\ell\right)})=w^{T}\cdot T_{\left(k,\ell\right)}$ where $w\in\mathbb{R}^{q\times q}$ is a weight matrix and $T_{\left(k,\ell\right)}$ is the translation map restricted to a $q\times q$ window. Equation 1 is applied to yield the full expression in Equation 2. In CNNs, the weight matrix is implicitly determined and optimized in the kernel matrix $\mathbf{K}$ via backpropagation during model training. The kernel performs feature extraction on the vectorized data to optimally search for the relevant topological signatures which it uses to distinguish between the tumor classes. We hope to use a similar approach in the future to adaptively optimize a weighting function in the vectorization operation.
@@ 4.1 Experimental Setup
We study three distinct representations of a histopathology slide: a grayscaled, thresholded image, global persistent homology computed for the entire image, and local persistent homology in the form of PHCs. Our objective is to compare the performance of models trained with each of these three representations with respect to multiple performance metrics. We also compare model performance of PHCs based on two different filtrations.
@@ 4.1.1 Histopathology Dataset
We apply CNNs to a dataset of histopathology slides for the diagnosis of Osteosarcoma (Ost.), a rare form of bone cancer. It is available at the Cancer Imaging Archive (Clark et al., 2013; Leavey et al., 2019). The dataset consists of 1144 RGB images with resolution (1024, 1024, 3). The images are separated into three classes: non-tumorous (47%), non-viable (or necrotic) tumor (23%), and viable tumor (30%). It is desirable that the dataset is balanced in each class with roughly 381 images (roughly a third of the total number of images). To achieve this, we resample 381 non-tumorous images and perform image augmentation (rotations and reflections) on the remaining classes to balance them. Since homology is invariant to rotations and reflections, the global PH data will contain multiple identical summaries. This illustrates that certain forms of data augmentation cannot be used when persistence is computed globally without introducing redundancies. The resulting experimental dataset contains 1143 RGB images and serves as our base-line comparison against the PH-based data.
@@ 4.1.2 PHC Data Generation
There are several preprocessing steps that are taken to condition the image data before computing persistent homology. These standard methods are applied to enhance the geometric features in the tissue structure, enabling more accurate and robust summaries. Note that we apply the same preprocessing to the data used for each of the three methods (that is: global persistence, local persistence) to ensure a fair comparison. The basic procedure is as follows:
@@
    The image is conditioned (truncated/thresholded) and the tissue architecture is emphasized using erosion.
@@
    All remaining pixels (the boundary of the cells) are appended to a data structure known as a simplex tree along with edges from neighboring pixels whose values are nonzero.
@@
Figure 6: Our preprocessing pipeline for image classification using PHC. As the convolution operator slides across the image several sub-tasks are performed. The sub-image is first conditioned before being used for complex creation. Next, the persistence of the complex is computed and vectorized. The data is collated and used for model training.
@@
The gray scale images are then conditioned used one of two techniques: thresholding or truncating. Thresholding is applied when computing the alpha complex, and truncation is applied for the other filtrations. For truncation, all pixel values above a parameter $\rho$ are retained, including their pixel intensity value. Both methods segment the tissue architecture with the former resulting in a binary image. This removes extraneous features, such as small, noisy holes between cells and within cells. We experimented with multiple truncation/threshold values ranging from 180-200 and found $k=200$ yielded the best performance. We also found that performance was worse if no truncation or thresholding was applied.
Erosion is applied once to “thicken” the boundary of the cells and fill in any small, “noisy” holes using a $2\times 2$ kernel.
We compute the persistent homology using multiple filtration methods. Please see Appendix Acknowledgments for the full list of results. We restrict our focus here to the two best performing filtrations: the alpha complex and a modified version of lower star filtration. The alpha complex filtration is computed by inputting the centers of the pixels which remain after thresholding into the implementation in GUDHI (The GUDHI Project, 2020). The second is a modified form of extended persistent homology on the lower star filtration. Here, the conditioned image data from Step 2 is converted into a data structure known as simplex tree, a trie data structure used to efficiently represent any general simplicial complex (Boissonnat and Maria, 2014). We create a simplex tree by adding vertices for each pixel, and edges between adjacent pixels in the truncated set. Each vertex is given a filtration value that correspond to the grayscale value of the pixel. For edges, its filtration value is determined by the max of its two vertices. This is called the adjacency complex of the image. The extended persistent homology of this filtration yields the same information as the birth times of the persistent homology intervals of the one described in Propositions 2 and 3. We found that this yielded better performance than the extended persistent homology of the lower star filtration.
@@
so $n=20$ was fixed.
The output PHC data of a single image is a 3-dimensional array of the form $(256,400,1)$ which corresponds to $256$ persistence images, one at each locality, with resolution $20\times 20$. Convolutions act on the array by comparing the entries of each topological summary from a particular locality with those of its neighbors. This enables the model to learn spatial similarities and multiscale patterns in the topology as the receptive field increases in subsequent convolution layers.
All persistence-based computations are performed using the computational topology library GUDHI (The GUDHI Project, 2020). Our pipeline is visually summarized in Figure 6. In its current form, most of the PHC data generation is a separate preprocessing step applied on the image data outside of the training cycle except for the optimizing the weights of $\mathbf{K}$. Further optimization can be performed on the parameter space which we plan to expand upon in future work, but is outside the scope of this study.
4.2 Intrinsic Dimensionality of Data
The manifold hypothesis posits that many high-dimensional datasets are supported on low-dimensional latent manifolds (Cayton and others, 2005). We hypothesize that PHCs reduce the intrinsic dimension of a dataset while retaining important geometric information.
That is, the persistent images are described by significantly fewer latent variables than the original image patches. There are multiple reasons why this might be beneficial:
  • 1.
    Increased memory and resource efficiency, and reduced number of parameters required during training.
  • 2.
    Lower intrinsic dimension is often correlated with better generalization from training to testing sets (Pope et al., 2021).
  • 3.
    The sample complexity to learn the data manifold depends exponentially on the intrinsic dimension. (Narayanan and Mitter, 2010).
A commonly used method to estimate the intrinsic dimension of each data representation is the Maximum Likelihood Estimator (MLE) proposed by (Levina and Bickel, 2004). The intrinsic dimension is estimated statistically by relating the density of points to the surface area of a spheres of varying dimensions. A complete explanation along with the mathematical construction of these estimators can be found in the references (Levina and Bickel, 2004; MacKay, David and Ghahramani, Zoubin, 2005).
@@ 4.3 Model, Training, and Implementation Details
We describe the CNN architecture used for our computational study. Further, our study indicates that a small CNN is sufficient to achieve high performance, consisting only of two convolution and pooling layers (which update the weights of $\mathbf{K}$) followed by three dense layers for the multilayer perceptron. These models are small enough that most modern laptops have sufficient hardware to train them.
All models are created and trained using TensorFlow (Abadi et al., 2016). The exact specifications of the architecture is as follows (see (Goodfellow et al., 2016) for an explanation of the terms used below). The kernel size is fixed to $(3,3)$ for both convolutions layers with a stride of two. ReLu is the default activation function used for all layers of the CNN except the last which uses the SoftMax function. The weights of the model are initialized using He Normal Initialization (He et al., 2015). We employ Adam (Kingma and Ba, 2015) for optimization with a fixed learning rate of $\alpha=0.001$. The loss functions used is categorical cross entropy given by the formula
@@
Similarly the convolution filter sizes vary from $\{8,16,32\}$ and layer size from $\{64,128,256,512,1028\}$. All of these hyperparameters were chosen using a Bayesian hyperparameter sweep implemented using Weights and Biases (Biewald, 2020) to maximize for accuracy.
In addition to comparing model performance between the three distinct representations, we also test the combination of image data with persistence (images + global PH or images + local PH). To train a model with the combined representations of image and PH data the CNN model architecture is slightly modified. Instead, it includes two separate feature extraction blocks with the same number of convolution and pooling layers as before using the same sweep configuration listed above. The image and PH data are fed through each separate block and concatenated together before being fed through the dense layers. We hypothesize that the inductive biases (translational invariance, locality, and hierarchical learning (Maurício et al., 2023)) that are built into CNNs enable the model to compare topological signatures in the persistence images across localities to find spatial relationships of local topologies. All five distinct training sets were trained over a 1000 model initializations.
@@ 4.4 Evaluation Metrics
$\textbf{Accuracy}=\frac{\text{Correct Classifications}}{\text{All Classifications}}.$